# Qizhe Zhang1,2  Aosong Cheng1  Ming Lu3  Zhiyong Zhuo1  Minqi Wang1 Jiajun Cao1  Shaobo Guo2  Qi She2  Shanghang Zhang1

### v1

@@ 5.3 FasterVLM with higher resolution
Some works [37, 33, 59] attempt to improve VLM performance on visual question answering tasks by increasing the input image resolution. However, this resolution increase introduces more visual tokens, further intensifying the computational load on the VLM. In this section, we apply FasterVLM to LLaVA-NeXT-7B, which can handle up to 2880 visual tokens. The results are presented in Tab. 3. Compared to LLaVA-1.5, LLaVA-NeXT involves a greater number of visual tokens, implying a higher degree of redundancy. At a 95% reduction ratio, retaining only 145 visual tokens, FasterVLM preserves 88.35% of the original performance, significantly outperforming FastV and FitPrune. This demonstrates the value of FasterVLM for high-resolution visual inputs.
5.4 FasterVLM with video understanding
@@
Table 4: Video understanding performance with Video-LLaVA across 4 commonly used video question answering benchmarks at reduction ratios ranging from 75% to 95%. The best average performance at each reduction ratio is shown in bold. Performance is evaluated using the first 1,000 samples from each benchmark and gpt-3.5-turbo is used for scoring.
Method Reduction Ratio # Token TGIF-QA MSVD-QA MSRVTT-QA ActivityNet-QA Average
Acc. Score Acc. Score Acc. Score Acc. Score Acc. Score
Video-LLaVA 0% 2048 19.80 2.53 70.50 3.93 57.50 3.50 43.60 3.81 100.00% 100.00%
FastV 75% 512 19.40 2.51 69.10 3.91 54.60 3.43 41.50 3.80 96.53% 99.12%
FasterVLM 17.90 2.48 70.20 3.95 56.60 3.50 44.70 3.86 97.73% 99.88%
FastV 90% 208 13.80 2.40 68.80 3.90 52.90 3.40 41.30 3.79 88.50% 97.63%
FasterVLM 15.60 2.39 69.10 3.92 55.30 3.43 44.50 3.82 93.76% 98.04%
FastV 95% 104 10.60 2.29 64.10 3.78 52.40 3.39 40.30 3.78 82.00% 95.64%
FasterVLM 14.00 2.34 65.30 3.79 53.80 3.40 43.70 3.79 89.28% 96.37%
@@
In addition to high-resolution images, video is another scenario with high redundancy in visual tokens. Thanks to its simple design, we can easily apply FasterVLM to Video-LLaVA, which accepts videos as input. Following [39] and [30], we conduct experiments on four video question answering benchmarks, using ChatGPT-Assistant for evaluation. Due to the commercial API usage limits, we follow [8] to use the first 1K samples from each benchmark in our experiments. The evaluation results are shown in Tab. 4. Video-LaVA processes 8 frames of 224-resolution video, totaling 2048 visual tokens. FasterVLM maintains 93.76% of the original performance at a 90% reduction ratio (retaining 130 tokens) and 89.28% with only 65 tokens (95% reduction ratio), significantly outperforming FastV. This improvement is due to the higher redundancy in temporally continuous video frames compared to single images, and the [CLS] attention accurately identifies key tokens within the video sequence, enabling FasterVLM to maintain strong performance even at high reduction ratios.
@@ 5.5 Ablation study
Table 5: Ablation studies of different pruning strategies on 3 benchmarks.
Pruning Strategy TextVQA POPE MME
LLaVA-1.5-7B 58.21 85.88 1506.47
(a) w/ randomly prune 90% before LLM 48.45 74.77 1227.07
(b) w/ prune 90% by patch attn before LLM 54.15 70.39 1278.05
(c) w/ prune 90% by [CLS] attn before LLM 55.28 75.85 1348.63
(d) w/ prune 90% by [CLS] attn at LLM layer 2 55.18 73.51 1338.43
(e) w/ prune 90% by [CLS] attn and merge 49.72 68.19 1207.31
Although our design is simple, it is both effective and efficient. To demonstrate this, we conduct a thorough ablation study on different strategies, presented in Tab. 5. Lines (a)-(c) examine different criteria for pruning visual tokens before the LLM, including random pruning (a), pruning based on the average self-attention among all image patches (b), and pruning based on [CLS] attention (c, used in our FasterVLM). Using image patch attention is less effective than [CLS] attention, as the [CLS] token contains more global information. Interestingly, random pruning does not perform poorly, further highlighting the redundancy of visual tokens in VLMs. In line (d), we prune tokens at the same layer as FastV but determine token importance based on [CLS] attention. This approach performs similarly to pruning before the LLM but results in lower inference efficiency. In line (e), we apply a commonly used strategy in token compression for VLMs, token merging, where each retained token is combined with its $k$ most similar tokens based on key similarity after selecting important tokens using [CLS] attention. Unfortunately, this strategy proves unsuitable for training-free token pruning, as greater merging led to significant performance degradation. Thus, FasterVLM adopts the simplest yet most effective method, directly pruning visual tokens before the LLM using [CLS] attention.
@@ 5.6 Efficiency analysis
Table 6: Inference efficiency comparison between FastV and FasterVLM with LLaVA-NeXT.
Method Reduction # Token FLOPs (T) Storage (MB) GPU Memory (GB) Throughput (it/s)
LLaVA-NeXT-7B 0% 2880 41.65 1440 17.93 2.40
FastV 50% 1440 21.11 765 17.98 3.22
FasterVLM 19.74 720 16.23 3.48
FastV 95% 145 6.14 225 17.98 4.56
FasterVLM 3.77 144 15.28 5.25
To demonstrate the efficiency of our FasterVLM, we perform a comparative analysis of FLOPs, cache storage, CUDA memory, and inference throughput with FastV on LLaVA-NeXT-7B. To get rid of the impact of output sequence length on the decoding time, we choose the SQA-IMG benchmark for analysis, where the model only outputs a single option. All experiments are performed on a single NVIDIA A100-80GB GPU. As shown in Tab. 6, at a 95% reduction ratio, FasterVLM achieves a ×$\times 2.2$ increase in throughput for LLaVA-NeXT. Compared to FastV with the same reduction ratio, FasterVLM requires less memory and achieves faster inference. This is because pruning within the LLM results in varying KV cache lengths, which is incompatible with current hardware. Moreover, the design of pruning before LLM enables compatibility with faster attention implementations like FlashAttention [10], which are infeasible for methods like FastV that require access to internal LLM attention information. Additional experiments are provided in the appendix.
6 Conclusion
In conclusion, we first reveal the inaccuracies in recent popular text-visual attention of VLMs for visual token pruning caused by the phenomena of attention shift and attention dispersion. Then, we introduce a straightforward and effective training-free visual token pruning method called FasterVLM, which evaluates the importance of visual tokens more accurately by the attentions between the [CLS] and image tokens from the visual encoder. By pruning the visual tokens before the LLM, we achieve better performance and faster inference than existing text-visual attention-based pruning methods.
Appendix
In this appendix, we first provide the details of the experimental setup in Appendix A, including information about the datasets, model architectures, and comparison methods. Then, in Appendix B, we offer a more detailed analysis of the attention in VLMs, along with more visualizations and insights. In Appendix C, we present additional experimental results, including results with richer details across different model architectures and visualizations of the ablation studies. Finally, Appendix D includes further efficiency analysis with FlashAttention.
Appendix A Details of Experimental Setup
A.1 Datasets
We evaluate our method on a total of 14 widely used benchmarks, including 10 image benchmarks and 4 video benchmarks. Each task is described as follows.
A.1.1 Image Benchmarks
@@
VQAv2 [15]. The VQAv2 benchmark evaluates the model’s visual recognition capabilities through open-ended questions. It consists of 265,016 images from MSCOCO dataset [31], with each image containing at least 3 questions. The dataset incorporates adversarially balanced question design, ensuring that each question corresponds to at least two images with completely different answers, preventing models from relying solely on statistical patterns to derive answers. We utilize the LLaVA subset of test-dev set for evaluation, which includes 107,394 image-question pairs. Each question is associated with 10 ground truth answers, and automatic evaluation metrics are used for scoring.
GQA [20]. The GQA benchmark focuses on evaluating structured understanding and reasoning abilities for scenes depicted in images. In addition to images and questions, it provides scene graph annotations derived from the Visual Genome dataset [24] for each image, including structured descriptions of objects, attributes, and their relationships within the scene. The questions are generated using the scene graphs, ensuring that each question corresponds to a clear semantic path. We use the accuracy on the test-dev set for evaluation, which contains 12,578 image-question pairs.
@@
TextVQA [50]. The TextVQA benchmark is designed to evaluate the model’s ability to recognize textual information within images, emphasizing the integration of optical character recognition (OCR) and natural language understanding. The images are primarily sourced from the Open Images v3 dataset [23] and contain a variety of scenarios such as signs, billboards, and product packaging that contain rich text information. In addition to raw images, reference OCR tokens are also provided. Answers to the questions may be directly derived from the text in the images or require contextual reasoning. We evaluate the performance on a validation set containing 5,000 image-question pairs.
@@ A.1.2 Video Benchmarks
ActivityNet-QA [63]. The ActivityNet-QA benchmark is based on the ActivityNet dataset [4], consisting of 5,800 complex web videos with an average length of 180 seconds. The benchmark includes 58K question-answer pairs, all of which are manually annotated to ensure higher quality.
A.2 Model Architectures
@@
A.3 Comparison Methods
FastV [8]. In vision-language models, visual tokens contribute significantly to computational overhead. FastV first identifies the inefficient visual attention phenomena in VLMs and proposes a straightforward method to address this. Specifically, it prunes visual tokens with the lowest Rvisual-text attention scores after layer 2 of the LLM, according to specific computation budget constraints, thereby accelerating VLM inference.
FitPrune [61]. Based on FastV, FitPrune investigates when and how to prune visual tokens more effectively. It uses binary search to keep the distribution of visual-text attention as unchanged as possible under given computational constraints, resulting in an optimal pruning strategy. Then, visual tokens are pruned according to the pre-calculated strategy, maximizing the preservation of original performance.
SparseVLM [65]. Unlike previous approaches, SparseVLM focuses on the influence of text prompts on visual-text attention. It proposes that not all text information is beneficial for pruning visual tokens. SparseVLM first selects text tokens relevant to the visual input and uses these raters to guide the pruning of visual tokens. Additionally, it employs a token recycle mechanism to retain more critical information, achieving notable performance improvements.
Appendix B Detailed Analysis of Attention in VLMs
B.1 Attention distribution
@@
B.2 Attention intensity
Figure 9: Visualizations of attention maps from the [CLS] token in the visual encoder and the last token in the language model. The attention in the visual encoder is concentrated, with higher attention to key foreground objects and regions that aggregate global information, while other areas receive little attention. In contrast, the attention in the language model is more dispersed and contains more noise, with important visual tokens receiving less attention. The highest attention tends to focus on the padding region, and attention is stronger in the lower part of the image than in the upper part.
@@ B.2 Attention intensity
Appendix C Additional Experimental Results
We provide further results for different methods on LLaVA-1.5-7B, LLaVA-1.5-13B, LLaVA-NeXT-7B, and Video-LLaVA in Tab. 7, Tab. 8, Tab. 9, and Tab. 10. Notably, the pruning scheme fitting algorithm of FitPrune is not compatible with the larger 13B language model or LLaVA-NeXT with dynamic number of visual tokens. Additionally, SparseVLM cannot reduce the number of visual tokens below 95% due to its pruned token recycling design. Our method achieves outstanding performance across all model architectures and different reduction ratios, demonstrating its simplicity and effectiveness. Notably, compared to the 7B version of LLaVA-1.5, increasing the size of the language model improves the performance of VLM on multi-modal tasks. Further improvements can be achieved with higher resolution image inputs. With just the 7B language model, LLaVA-NeXT outperforms LLaVA-1.5, which uses the 13B language model. Additionally, video inputs, which contain more redundancy than single images, benefit more from visual token pruning.
Fig. 11 visualizes the performance of FasterVLM on LLaVA-1.5-7B across different reduction ratios. Our method maintains strong performance even at a 95% reduction ratio, significantly outperforming versions that do not use visual input. This demonstrates the critical importance of a small number of key visual tokens in VLM performance. Fig. 11 presents the results of ablation experiments on various pruning strategies. By performing visual token pruning based on the [CLS] attention before LLM, our FasterVLM achieves remarkable efficiency. This design, while simple, proves to be highly effective in preserving model performance while reducing computational complexity.
Table 7: Performance comparison of different methods with LLaVA-1.5-7B across 10 benchmarks under varying reduction ratios. For each reduction ratio, the best average performance is shown in bold, and the second best one is underlined. 100% reduction ratio represents using no input images.
Method Reduction Ratio # Token VQAv2 GQA VizWiz SQA-IMG TextVQA POPE MME MMB MMB-CN MM-Vet Average
LLaVA-1.5-7B 0% 576 78.52 61.94 50.06 69.51 58.21 85.88 1506.47 64.69 58.08 31.30 100.00%
FastV 25% 432 78.45 61.69 50.30 69.31 58.09 85.29 1528.16 64.52 58.76 32.80 100.60%
FitPrune 78.49 61.93 50.07 69.51 58.27 85.94 1512.72 64.60 58.33 31.50 100.15%
SparseVLM 78.13 61.38 50.38 68.62 57.44 84.56 1475.47 64.78 57.30 32.40 99.54%
FasterVLM 78.39 61.51 50.01 68.57 57.91 85.92 1501.10 65.12 58.51 32.40 100.18%
FastV 50% 288 77.67 60.05 50.53 68.96 58.25 82.45 1513.06 64.26 58.16 31.70 99.33%
FitPrune 78.41 61.70 50.04 69.16 58.26 85.37 1499.70 64.60 58.16 31.10 99.73%
SparseVLM 76.67 58.78 51.03 68.57 57.49 83.87 1458.79 63.14 56.87 31.50 98.52%
FasterVLM 77.86 60.64 50.45 68.37 57.90 86.20 1471.96 63.83 56.70 34.80 100.12%
FastV 75% 144 74.07 56.58 51.29 69.11 57.38 73.74 1463.39 64.00 57.22 28.60 95.80%
FitPrune 76.14 59.38 51.30 69.01 56.49 80.75 1472.86 63.92 57.65 28.40 97.22%
SparseVLM 72.76 55.11 51.46 69.36 55.99 77.57 1351.65 59.54 51.03 29.90 93.84%
FasterVLM 76.19 58.34 51.97 67.92 57.07 83.46 1433.76 62.54 57.13 34.20 98.75%
FastV 90% 58 65.38 51.20 51.84 69.81 54.75 57.30 1210.36 59.97 51.72 27.20 87.97%
FitPrune 62.76 49.96 50.85 68.22 50.35 53.81 1147.46 56.27 45.53 21.80 82.07%
SparseVLM 62.90 48.86 49.36 67.23 48.99 65.82 1030.61 49.05 35.40 18.60 78.13%
FasterVLM 71.92 54.91 53.01 68.91 55.28 75.85 1348.63 60.57 54.90 30.10 94.24%
FastV 95% 29 55.92 46.03 49.10 70.00 51.56 35.47 971.56 50.17 42.18 18.90 74.93%
FitPrune 52.39 43.60 48.61 68.32 46.75 31.17 855.21 39.69 29.98 18.00 67.64%
FasterVLM 66.75 51.51 52.67 69.56 53.09 67.24 1254.80 58.51 51.98 27.50 89.41%
LLaVA-1.5-7B 100% 0 40.73 37.38 45.36 63.06 41.43 47.21 719.10 20.02 17.70 11.30 56.50%
Table 8: Performance comparison of different methods with LLaVA-1.5-13B across 10 benchmarks under varying reduction ratios. For each reduction ratio, the best average performance is shown in bold, and the second best one is underlined. 100% reduction ratio represents using no input images.
Method Reduction Ratio # Token VQAv2 GQA VizWiz SQA-IMG TextVQA POPE MME MMB MMB-CN MM-Vet Average
LLaVA-1.5-13B 0% 576 80.00 63.25 53.61 72.78 61.19 85.99 1531.19 68.47 63.49 36.30 100.00%
FastV 25% 432 79.96 63.11 53.79 72.78 61.29 85.86 1543.97 68.38 63.40 35.20 99.76%
SparseVLM 79.25 61.29 53.60 73.77 60.46 72.45 1481.40 68.30 61.77 35.50 97.20%
FasterVLM 79.61 61.24 53.16 72.98 60.35 86.73 1487.34 67.61 63.40 36.70 99.21%
FastV 50% 288 79.54 62.59 54.34 73.13 60.86 85.15 1545.14 68.47 63.23 34.80 99.51%
SparseVLM 78.48 59.90 53.08 74.02 59.48 71.30 1497.39 66.67 61.94 36.50 96.69%
FasterVLM 79.03 61.01 52.65 73.62 59.99 86.05 1530.41 67.70 62.80 36.80 99.18%
FastV 75% 144 77.24 59.87 54.82 74.02 60.07 79.43 1493.51 67.27 62.63 33.20 97.16%
SparseVLM 76.06 57.97 53.13 73.67 57.94 68.61 1499.49 64.52 59.11 35.00 94.32%
FasterVLM 77.36 58.74 52.74 73.48 58.99 83.10 1467.00 67.10 62.54 36.30 97.43%
FastV 90% 58 70.27 54.92 54.78 72.43 55.64 67.26 1359.69 63.83 59.71 29.40 90.26%
SparseVLM 68.27 54.43 50.45 70.35 52.56 62.63 1285.26 58.16 54.30 27.20 85.02%
FasterVLM 73.08 55.98 54.00 73.72 57.35 74.71 1370.77 65.21 61.08 33.90 93.68%
FastV 95% 29 62.25 50.34 52.96 73.18 52.08 49.83 1165.70 56.44 51.29 24.00 80.53%
FasterVLM 67.85 52.62 53.11 72.83 54.82 65.90 1267.09 62.11 56.87 31.60 88.35%
LLaVA-1.5-13B 100% 0 41.40 38.40 45.15 66.63 43.51 1.45 622.91 22.77 18.13 13.20 49.99%
Table 9: Performance comparison of different methods with LLaVA-NeXT-7B across 10 benchmarks under varying reduction ratios. For each reduction ratio, the best average performance is shown in bold, and the second best one is underlined. 100% reduction ratio represents using no input images.
Method Reduction Ratio # Token VQAv2 GQA VizWiz SQA-IMG TextVQA POPE MME MMB MMB-CN MM-Vet Average
LLaVA-NeXT-7B 0% 2880 81.21 62.93 55.21 69.66 59.59 86.32 1513.78 67.70 58.85 42.60 100.00%
FastV 25% 2160 81.12 62.50 55.05 69.31 59.69 86.27 1506.28 67.61 59.02 41.70 99.61%
SparseVLM 81.14 62.55 55.20 68.47 60.26 73.15 1507.75 66.07 58.59 41.90 97.86%
FasterVLM 81.18 62.81 56.12 70.70 59.65 86.26 1492.20 67.35 58.33 44.30 100.41%
FastV 50% 1440 80.71 61.76 54.89 69.06 59.55 85.46 1490.34 67.35 58.51 41.20 98.91%
SparseVLM 80.92 62.04 55.71 68.07 60.00 73.42 1484.92 65.72 58.85 39.90 97.14%
FasterVLM 80.72 62.66 55.82 69.41 59.72 86.66 1521.87 67.87 58.85 44.40 100.53%
FastV 75% 720 78.90 60.38 54.22 69.81 58.39 83.09 1477.31 65.64 57.04 41.10 97.37%
SparseVLM 78.86 60.88 55.55 67.48 58.08 70.99 1446.10 63.83 57.04 38.00 94.95%
FasterVLM 79.25 61.31 56.24 68.82 59.33 85.50 1480.68 67.53 59.19 40.40 98.73%
FastV 90% 288 71.94 55.86 53.07 69.26 55.69 71.66 1282.86 61.60 51.89 33.70 89.24%
SparseVLM 71.62 56.12 53.16 68.62 51.97 63.23 1332.22 54.47 50.69 24.70 84.52%
FasterVLM 75.21 58.12 56.92 68.12 57.57 80.00 1370.11 63.32 54.47 35.70 93.55%
FastV 95% 144 61.84 49.83 51.25 68.52 51.85 51.66 1079.46 54.90 45.36 21.90 77.43%
FasterVLM 70.63 54.73 56.27 68.86 55.97 72.89 1225.96 60.48 53.09 31.90 88.85%
LLaVA-NeXT-7B 100% 0 40.59 37.93 46.01 64.01 37.57 23.40 601.93 21.05 17.70 13.10 50.73%
Table 10: Performance comparison of different methods with Video-LLaVA across 4 video question answering benchmarks under varying reduction ratios. Evaluation was performed using the first 1,000 samples from each benchmark with gpt-3.5-turbo assistant. For each reduction ratio, the best average performance is highlighted in bold, and 100% reduction ratio represents using no video frames.
Method Reduction Ratio # Token TGIF-QA MSVD-QA MSRVTT-QA ActivityNet-QA Average
Acc. Score Acc. Score Acc. Score Acc. Score Acc. Score
Video-LLaVA 0% 2048 19.80 2.53 70.50 3.93 57.50 3.50 43.60 3.81 100.00% 100.00%
FastV 25% 1536 19.60 2.54 70.50 3.94 55.40 3.47 44.20 3.86 99.18% 100.24%
FasterVLM 19.80 2.53 72.00 3.96 57.10 3.48 45.00 3.88 101.16% 100.45%
FastV 50% 1024 19.50 2.54 70.40 3.94 54.80 3.46 43.40 3.81 98.30% 99.79%
FasterVLM 19.20 2.51 71.90 3.96 56.70 3.47 44.90 3.87 100.14% 100.11%
FastV 75% 512 19.40 2.51 69.10 3.91 54.60 3.43 41.50 3.80 96.53% 99.12%
FasterVLM 17.90 2.48 70.20 3.95 56.60 3.50 44.70 3.86 97.73% 99.88%
FastV 90% 208 13.80 2.40 68.80 3.90 52.90 3.40 41.30 3.79 88.50% 97.63%
FasterVLM 15.60 2.39 69.10 3.92 55.30 3.43 44.50 3.82 93.76% 98.04%
FastV 95% 104 10.60 2.29 64.10 3.78 52.40 3.39 40.30 3.78 82.00% 95.64%
FasterVLM 14.00 2.34 65.30 3.79 53.80 3.40 43.70 3.79 89.28% 96.37%
Video-LLaVA 100% 0 5.70 1.96 23.40 1.77 24.50 1.87 39.10 3.66 48.57% 68.01%
Figure 10: Performance of FasterVLM on LLaVA-1.5-7B across 10 benchmarks at varying reduction ratios from 25% to 100%.
Figure 11: Ablation study on different pruning strategies in FasterVLM, showing performance across 10 benchmarks.
Appendix D Efficiency Analysis with FlashAttention
In Tab. 11, Tab. 12 and Tab. 13, we compare the computational efficiency between FastV and our FasterVLM under LLaVA-1.5-7B, LLaVA-1.5-13B, and LLaVA-NeXT-7B, respectively. Unlike FastV, which prune visual token within the LLM, FasterVLM prunes tokens before the LLM, enabling compatibility with FlashAttention. This design results in significantly higher efficiency. Note that the original implementation of SDPA also includes FlashAttention, so its computational efficiency is comparable to that of FlashAttention2, with only slight differences. All efficiency analyses are performed on a single NVIDIA A100-80GB GPU, evaluated using the POPE benchmark.
Table 11: Efficiency comparison between FastV and FasterVLM under LLaVA-1.5-7B.

### v2

@@ 5.3 VisPruner with higher resolution
Some works [39, 35, 58] attempt to improve VLM performance on visual question answering tasks by increasing the input image resolution. However, this resolution increase introduces much more visual tokens, further intensifying the computational load on the VLM. In this section, we apply VisPruner to LLaVA-NeXT-7B, which can handle up to 2880 visual tokens. The results are presented in Tab. 2. Compared to LLaVA-1.5, LLaVA-NeXT involves a greater number of visual tokens, implying a higher degree of redundancy. At a 94.4% reduction ratio, retaining only 160 visual tokens, VisPruner preserves 86.7% of the original performance, significantly outperforming FastV (74.7%), SparseVLM (74.5%) as well as VisionZip (82.3%), demonstrating the value of VisPruner for high-resolution visual inputs.
5.4 VisPruner with video understanding
Method TGIF-QA MSVD-QA MSRVTT-QA Average
Acc. Score Acc. Score Acc. Score Acc. Score
Upper Bound, All 2048 Tokens ($\mathbf{100\%}$)
Video-LLaVA 19.8 2.53 70.5 3.93 57.5 3.50 49.3 3.32
Retain 455 Tokens ($\downarrow\mathbf{77.8\%}$)
FastV [9](ECCV24) 19.2 2.50 69.1 3.91 54.4 3.42 47.6 3.28
VisPruner(Ours) 18.4 2.49 70.2 3.95 56.7 3.50 48.4 3.31
Retain 227 Tokens ($\downarrow\mathbf{88.9\%}$)
FastV [9](ECCV24) 14.3 2.42 68.9 3.90 53.0 3.40 45.4 3.24
VisPruner(Ours) 15.9 2.41 69.3 3.92 55.6 3.45 46.9 3.26
Retain 114 Tokens ($\downarrow\mathbf{94.4\%}$)
FastV [9](ECCV24) 10.6 2.29 64.1 3.78 52.4 3.39 42.4 3.15
VisPruner(Ours) 14.1 2.35 65.4 3.79 54.1 3.41 44.5 3.18
@@ 5.4 VisPruner with video understanding
Table 3: Video understanding performance with Video-LLaVA across three commonly used video question answering benchmarks. Performance is evaluated using the first 1,000 samples from each benchmark and gpt-3.5-turbo is used for scoring.
@@
Video understanding is another scenario with high visual redundancy. We also apply our VisPruner to Video-LLaVA, which accepts videos as input. Following [40] and [32], we conduct experiments on three video question answering benchmarks, using ChatGPT-Assistant for evaluation. Due to the commercial API usage limits, we follow [9] to use the first 1K samples from each benchmark in our experiments. The evaluation results are shown in Tab. 3. Video-LLaVA processes 8 frames of 224-resolution video, totaling 2048 visual tokens. VisPruner maintains 95.1% of the original performance at a 88.9% reduction ratio (retaining 227 tokens) and 90.3% with only 114 tokens (94.4% reduction ratio), significantly outperforming FastV (86%). This improvement is due to the higher redundancy in temporally continuous video frames compared to single images, and the [CLS] attention accurately identifies key tokens within the video sequence, enabling VisPruner to maintain strong performance even at high reduction ratios.
5.5 VisPruner with other VLM architecture
Method $\text{VQA}^{\text{V2}}$ GQA VizWiz $\text{SQA}^{\text{IMG}}$ $\text{VQA}^{\text{Text}}$ Acc. Rel.
Upper Bound, All 256 Tokens ($\mathbf{100\%}$)
Qwen-VL-7B 78.8 59.3 35.2 67.1 63.8 60.8 100.0%
Retain 128 Tokens ($\downarrow\mathbf{50\%}$)
FastV [9](ECCV24) 76.5 56.9 32.7 65.3 58.2 57.9 94.9%
VisPruner(Ours) 77.4 57.8 33.4 65.9 59.6 58.8 96.4%
Upper Bound, All 576 Tokens ($\mathbf{100\%}$)
InternVL-Chat-13B 79.3 62.9 52.6 66.3 57.0 63.6 100.0%
Retain 144 Tokens ($\downarrow\mathbf{75\%}$)
FastV [9](ECCV24) 74.1 58.2 50.8 66.6 55.6 61.1 96.1%
VisPruner(Ours) 76.7 60.2 51.9 67.3 55.3 62.3 97.9%
Upper Bound, All 1225 Tokens ($\mathbf{100\%}$)
CogVLM-chat-1.1-17B 80.9 58.2 49.6 68.4 69.2 65.3 100.0%
Retain 123 Tokens ($\downarrow\mathbf{90\%}$)
FastV [9](ECCV24) 74.2 40.3 42.9 63.5 41.9 52.6 80.2%
VisPruner(Ours) 74.6 48.4 46.6 68.2 59.6 59.5 91.0%
Table 4: Performance of VisPruner on different VLM architectures. Acc. denotes the average accuracy across benchmarks, Rel. represents the average percentage of performance maintained.
In addition to the LLaVA series, we also apply VisPruner to other model architectures, including the dominant Qwen-VL [3], InternVL [10], and CogVLM [54]. For Qwen-VL, which uses a Q-Former as a multimodal projector, we first extract attention for each patch token from the visual encoder, and then track each query’s attention towards patch tokens in the Q-Former. The integration of these two components yields the visual attention for the final query sequence. InternVL and CogVLM both use MLPs as projectors, so we apply VisPruner to them in the same manner as with LLaVA. The performance at different reduction ratios are shown in Tab. 4. Our VisPruner performs well across a variety of VLM architectures and reduction ratios, consistently outperforming the text-visual attention-based FastV, proving that visual cues can be exploited for effective token pruning across different architectures.
@@ 5.6 Ablation study
Figure 5: Ablation study of the core components. Random denotes randomly selecting tokens, Important refers to selecting only tokens with high attention scores, and VisPruner represents the final version of our method, which also includes a diverse set of tokens with low redundancy as a complement.
To verify the effectiveness of the two core components of our method, we conduct an ablation experiment. We retain 64 visual tokens from different sources in LLaVA-1.5-7B and observe their performance on the TextVQA and POPE benchmarks, as shown in Fig. 5. Randomly selecting tokens results in significant performance degradation due to the loss of important visual information. Selecting important tokens through visual attention led to a performance improvement of 5% and 1.5% on the two benchmarks. However, as attention in the visual encoder tends to focus on the foreground, retaining only important tokens results in the loss of information related to the background. Here, we also select a set of diverse tokens based on similarity from the remaining as a complement, which led to additional gains of 2.3% and 3.6%, respectively, demonstrating the importance of this information complementation for the model.
@@ 5.7 Efficiency analysis
Method # Token
FLOPs
Storage
GPU Memory
CUDA Time
LLaVA-NeXT-7B 2880 43.6 1440 17.0 313
FastV [9](ECCV24) 640 13.5 380 16.9 148
VisPruner(Ours) 11.5 360 14.8 117
FastV [9](ECCV24) 160 6.3 95 16.9 112
VisPruner(Ours) 3.8 80 14.7 78
Table 5: Efficiency analysis with LLaVA-NeXT-7B on POPE. At the same reduction ratio, VisPruner requires less memory and achieves faster inference compared to FastV.
To demonstrate the efficiency of VisPruner, we perform a comparative analysis of FLOPs, cache storage, GPU memory, and CUDA latency with FastV on LLaVA-NeXT-7B. All experiments are performed on a single NVIDIA A100-80GB GPU, evaluated using the POPE benchmark. As shown in Tab. 5, by retaining only 160 tokens, VisPruner achieves a ×$\mathbf{\times 4}$ reduction in inference latency for LLaVA-NeXT. Compared to FastV with the same reduction ratio, VisPruner requires less memory and achieves faster inference. Moreover, the design of pruning before the language model enables compatibility with faster attention implementations like FlashAttention [12], which are infeasible for methods like FastV that require access to the text-visual attention information with in the language model.
6 Discussion
In this discussion, we attempt to explain the outstanding performance of VisPruner. Despite the inherent redundancy in visual inputs, the distribution of visual information within an image is uneven. The semantic information in the foreground region is more dense, while is relatively sparse in the background area. Some concurrent works have also begun to utilize visual information to guide token pruning.  [53] leverages visual attention to select key tokens, resulting in the loss of background information. Other studies [45, 59] attempt to merge redundant tokens with key tokens, but the fusion process introduces more noise, leading to performance degradation without further training.
We propose a novel method that first uses attention to select important tokens, and then removes redundancy from the remaining ones based on similarity, ultimately retaining a set of diverse tokens as a supplement. This process can be viewed as dynamic pruning across areas with varying information densities; a higher proportion of tokens is retained in foreground areas, while fewer are kept in background areas. Based on this explanation, dynamically partitioning and finely pruning visual tokens according to varying information densities is an interesting direction for future work.
Supplementary Material
7 Details of experimental setup
7.1 Datasets
We evaluate our method on a total of 13 widely used benchmarks, including 10 image benchmarks and 3 video benchmarks. Each task is described as follows.
7.1.1 Image benchmarks
@@
VQAv2 [16]. The VQAv2 benchmark evaluates the model’s visual recognition capabilities through open-ended questions. It consists of 265,016 images from MSCOCO dataset [33], with each image containing at least 3 questions. The dataset incorporates adversarially balanced question design, ensuring that each question corresponds to at least two images with completely different answers, preventing models from relying solely on statistical patterns to derive answers. We utilize the test-dev set for evaluation, which includes 107,394 image-question pairs. Each question is associated with 10 ground truth answers, and automatic evaluation metrics are used for scoring.
GQA [22]. The GQA benchmark focuses on evaluating structured understanding and reasoning abilities for scenes depicted in images. In addition to images and questions, it provides scene graph annotations derived from the Visual Genome dataset [26] for each image, which include structured descriptions of objects, attributes, and their relationships within the scene. The questions are generated using the scene graphs and a pre-designed engine, ensuring that each question corresponds to a clear semantic path. We use the accuracy on the test-dev set for evaluation, which contains 12,578 image-question pairs.
@@
TextVQA [46]. The TextVQA benchmark is designed to evaluate the model’s ability to recognize textual information within images, emphasizing the integration of optical character recognition (OCR) and natural language understanding. The images are primarily sourced from the Open Images v3 dataset [25] and contain a variety of scenarios such as signs, billboards, and product packaging that contain rich text information. In addition to raw images, reference OCR tokens are also provided. Answers to the questions may be directly derived from the text in the images or require contextual reasoning. We evaluate the model’s performance on a validation set containing 5,000 image-question pairs.
@@ 7.1.2 Video benchmarks
7.2 Model architectures
@@
Qwen-VL [3]. Qwen-VL is another widely used open-source vision-language model. Similar to LLaVA, it includes a visual encoder (OpenCLIP) and a text decoder (Qwen LLM). For the vision-text connector, Qwen-VL employs a vision-language adapter, which transforms image inputs into fixed-length token sequences via cross-attention. After three stages of training, Qwen-VL achieves strong vision-language understanding capabilities. And Qwen-VL-Chat is further fine-tuned based on Qwen-VL to enhance its performance in conversational tasks.
@@
8 Detailed Analysis of Attention in VLMs
8.1 Attention distribution
@@
8.2 Attention intensity
Figure 9: Visualizations of attention maps from the [CLS] token in the visual encoder and the last token in the language model.
@@ 8.2 Attention intensity
9 Efficiency Analysis with FlashAttention
In Tab. 6, Tab. 7 and Tab. 8, we compare the computational efficiency between FastV and our VisPruner under LLaVA-1.5-7B, LLaVA-1.5-13B, and LLaVA-NeXT-7B, respectively. Unlike FastV, which prune visual token within the LLM, VisPruner prunes tokens before the LLM, enabling compatibility with FlashAttention. This design results in significantly higher efficiency. Note that the original implementation of SDPA also includes FlashAttention, so its computational efficiency is comparable to that of FlashAttention2, with only slight differences. All efficiency analyses are performed on a single NVIDIA A100-80GB GPU, evaluated using the POPE benchmark.